# Graph Classification Gaussian Processes via Hodgelet Spectral Features

### v1

@@ Abstract
The problem of classifying graphs is ubiquitous in machine learning. While it is standard to apply graph neural networks for such tasks, Gaussian processes can also be used, by transforming graph features into the spectral domain, and using the resulting spectral features as input points. However, this approach only takes into account features on vertices, whereas some graph data also support features on edges. In this work, we present a Gaussian process-based classification algorithm that can utilise vertex and/or edges features to help classify graphs. Furthermore, we take advantage of the Hodge decomposition of vertex and edge features to increase the flexibility of the model, which can be beneficial on some tasks.
@@ 1 Introduction
Classification is omnipresent in machine learning, yet typically assumes the data to be Euclidean. Extending this task to non-Euclidean domains, such as graphs, is challenging due to their irregularity, varying sizes, and multi-site information (e.g. vertices and edges). However, classifying graphs is of critical importance in scientific and industrial applications, being used for instance to predict properties of molecules, or discovering new drugs [1, 2]. Graph neural networks [3] are usually the model of choice for such applications, however, a downside is that they may require large datasets to train. Gaussian processes (GP) [4], on the other hand, prove to be a data-efficient and interpretable modelling choice, not requiring validation sets to tune hyperparameters, while providing robust uncertainty estimates on predictions. This makes them well-suited for a range of situations, such as in small-data regimes and high-risk decision-making tasks that require reliable uncertainty estimates.
@@
In a recent work, Opolka et al. [5] introduced a GP-based algorithm that is capable of classifying graphs. Their method relies on tools developed in the graph signal processing literature [6], such as the graph Fourier transform [7] and the spectral graph wavelet transform [8], to convert features in the graph domain into spectral features in the Euclidean domain. The latter can then be passed as input points to a GP, which is employed for classification via approximate inference [9]. More specifically, graph spectral methods provide a spectral representation of features on vertices. This representation is defined in the eigenspace of the graph Laplacian. Closely related are techniques developed in the early graph neural network literature, for instance Spectral Networks [10] and ChebNet [11], which rely on graph Fourier/wavelet transforms to define a notion of convolution on graphs.
While the method proposed in Opolka et al. [5] allows processing of features living on the vertices, it cannot straightforwardly take into account edge features. However, information about the edges of a graph can be as useful as the information on the vertices, for instance representing important quantities such as flows [12] and bond energies [13].
@@
Our paper aims to fill this gap by proposing a GP-based classification algorithm that can naturally incorporate features on vertices, edges and, more generally, simplices, building on recent work defining GPs on simplicial complexes [14] and cellular complexes [15]. Furthermore, we utilise the celebrated Hodge decomposition on vertex, edges and simplices features. This splits the aforementioned eigenspace into three canonical subspaces, called Hodge subspaces, each with distinct properties. This enables the modelling of each component separately, using different kernels. This idea have been used successfully in past works [16, 17, 18, 14], allowing for greater model flexibility. To the best of our knowledge, we are not aware that these recent extensions and techniques have been applied in the context of graph classification. We demonstrate empirically that these extensions achieve similar or better performance than the GP-based method of Opolka et al. [5].
Finally, while we only consider graph data here, our method can be extended easily to higher-order networks, such as simplicial complexes, cellular complexes, and hypergraphs, that can describe polyadic interaction [19, 20, 21, 22], generalising the dyadic interaction present in graphs (see Appendix C).
@@ 2 Gaussian Processes for Graph Classification
We assume a dataset containing $M$ undirected graphs $\mathcal{G}^{(1)},\dots,G^{(M)}$ and labels $y^{(i)}\in\mathbb{Z}$, for $i=1,\ldots,M$. We give each graph an orientation and emphasise that this choice is arbitrary. Let $\mathcal{V}^{(i)}$ be the vertex set and $\mathcal{E}^{(i)}$ be the edge set, both finite. Each graph $\mathcal{G}^{(i)}=(\mathcal{V}^{(i)},\mathcal{E}^{(i)})$ has $N_{v}^{(i)}$ vertices and $N_{e}^{(i)}$ edges, and an incidence matrix $\boldsymbol{B}_{ve}^{(i)}\in\mathbb{Z}^{N_{v}^{(i)}\times N_{e}^{(i)}}$. The latter encodes the incidence of each vertex with each edge, and furthermore defines the graph Laplacian $\boldsymbol{L}_{v}^{(i)}\coloneq\boldsymbol{B}_{ve}^{(i)}\boldsymbol{B}_{ve}^{(i)\top}\in\mathbb{Z}^{N_{v}^{(i)}\times N_{v}^{(i)}}$. By considering $3$-cliques (see Appendix C.2), we obtain the edge-to-triangle incidence matrix $\boldsymbol{B}_{et}^{(i)}\in\mathbb{Z}^{N_{e}^{(i)}\times N_{t}^{(i)}}$, where $N_{t}^{(i)}$ is the number of triangles in $\mathcal{G}^{(i)}$. Likewise, we can define the graph Helmholtzian $\boldsymbol{L}_{e}^{(i)}\coloneq\boldsymbol{B}_{ve}^{(i)\top}\boldsymbol{B}_{ve}^{(i)}+\boldsymbol{B}_{et}^{(i)}\boldsymbol{B}_{et}^{(i)\top}\in\mathbb{Z}^{N_{e}^{(i)}\times N_{e}^{(i)}}$, which applies to edges rather than vertices. Let $\mathcal{V}^{(i)}\rightarrow\mathbb{R}^{D_{v}}$ and $\mathcal{E}^{(i)}\rightarrow\mathbb{R}^{D_{e}}$ be two functions defined on vertices and edges, respectively, for $D_{v},D_{e}\in\mathbb{N}$. By introducing an ordering on vertices and edges, we can represent these functions trivially by the matrices $\mathbb{R}^{D_{v}\times N_{v}^{(i)}}$ and $\mathbb{R}^{D_{e}\times N_{e}^{(i)}}$, respectively. These are understood as vertex (and edge) feature matrices with $D_{v}$ (and $D_{e}$) channels. We denote vertex features of channel $1\leq d\leq D_{v}$ by the column vector $\boldsymbol{x}^{(i)}_{vd}\in\mathbb{R}^{N_{v}^{(i)}}$, and edge features of channel $1\leq d\leq D_{e}$ by the column vector $\boldsymbol{x}^{(i)}_{ed}\in\mathbb{R}^{N_{e}^{(i)}}$. Note that graphs may have vertex features, edge features, or both.
@@ 2.1 Wavelet transforms on graphs
The key idea behind our graph classification algorithm is to convert vertex and edge features in the graph domain into spectral features in the Euclidean domain, enabling the use of standard GP classification using these transformed features. To this end, we first consider the eigendecomposition of the graph Laplacian and Helmholtzian
@@
and define the graph Fourier transform as the projection of graph features onto the eigenbases,
@@
Vertex and edge Fourier coefficients, respectively, $\hat{\boldsymbol{x}}^{(i)}_{vd}$ and $\hat{\boldsymbol{x}}^{(i)}_{ed}$, are invariant to vertex and edge ordering, making them sound choices for constructing spectral features. They capture global properties of the original features. However, it is often beneficial to use spatially localised information, i.e. focusing only on local neighbourhoods, and use a multi-scale resolution. A way to achieve this is to use the graph wavelet transform [8] (see Appendix B). Wavelet functions, $b:\mathbb{R}\rightarrow\mathbb{R}$ and $d:\mathbb{R}\rightarrow\mathbb{R}$, operate as band-pass filters, and consist of scaled and shifted variants of a mother wavelet. Scaling functions, $a:\mathbb{R}\rightarrow\mathbb{R}$ and $c:\mathbb{R}\rightarrow\mathbb{R}$, are low-pass filters. This yield the functions on eigenvalues,
@@
where $L_{v},L_{e}$ are the number of scales, $W_{v},W_{e}$ the number of filters, and $\Theta_{vj},\Theta_{ej}$ are trainable parameters controlling the scales. Vertex and edge wavelet coefficients are given, respectively, by
@@ 2.2 Hodge decomposition
The Hodge decomposition [23] (see Appendix C.3) states that vertex and edge features can be canonically decomposed into the sum of three orthogonal parts: exact, co-exact, and harmonic. This leads to breaking down the eigenbases according to the Hodge subspaces,
@@
The co-exact eigenbasis $\boldsymbol{U}^{(i)}_{vc}$ corresponds to the non-zero eigenvectors of $\boldsymbol{L}_{v}^{(i)}$, and the harmonic eigenbasis $\boldsymbol{U}^{(i)}_{vh}$, the zero ones. The exact and co-exact eigenbases $\boldsymbol{U}^{(i)}_{ee}$ and $\boldsymbol{U}^{(i)}_{ec}$, respectively, are the non-zero eigenvectors of $\boldsymbol{B}_{ve}^{(i)\top}\boldsymbol{B}_{ve}^{(i)}$ and $\boldsymbol{B}_{et}^{(i)}\boldsymbol{B}_{et}^{(i)\top}$, respectively, and the harmonic eigenbasis $\boldsymbol{U}^{(i)}_{eh}$, the zero ones of $\boldsymbol{L}_{e}^{(i)}$.
@@ 2.3 Hodgelet spectral features
We generate spectral features from wavelet coefficients, and then use them in downstream GP classification tasks (see Appendix A). First, we take the wavelet transform (5) with respect to the decomposed eigenbasis $\boldsymbol{U}_{v}^{(i)}$ (6), yielding two wavelet coefficients, corresponding to the co-exact and harmonic parts. For every filter $1\leq j\leq W_{v}$, and every channel $1\leq d\leq D_{v}$, we concatenate the $2$-norm of the coefficient $\hat{\boldsymbol{x}}^{(i)}_{vdj}$. This results in two column vectors $\boldsymbol{v}^{(i)}_{c},\boldsymbol{v}^{(i)}_{h}\in\mathbb{R}^{W_{v}D_{v}}$, which we call vertex Hodgelet spectral features. We apply a similar procedure to get edge Hodgelet spectral features $\boldsymbol{e}^{(i)}_{e},\boldsymbol{e}^{(i)}_{c},\boldsymbol{e}^{(i)}_{h}\in\mathbb{R}^{W_{e}D_{e}}$. These are employed as input points for our additive Hodgelet kernel
@@
where each $\kappa_{\bullet\sqbullet}$ is a kernel, such as the squared exponential or Matérn kernel function. The $\Theta_{\bullet j}$ parameters are optimised jointly with the kernel hyperparameters.
@@ 3 Experiments
The aim of the experiments is two-fold: (1) we validate the added flexibility given by the Hodge decomposition, and (2) we demonstrate that when edge features are present, it is more beneficial to work with edges directly, rather than converting a graph to a line-graph.
The latter point has been observed in previous works, such as Schaub et al. [20] and Yang et al. [14], and our experiments further validate their conclusions.
@@ 3.1 Graph classification benchmarks
In our first experiment, we compare our method to Opolka et al. [5] on standard graph classification benchmark datasets that only contain vertex features. Specifically, we compare their algorithm, which we refer to the wavelet-transform GP (WT-GP), which does not use the Hodge decomposition, to our method, WT-GP-Hodge, which employs the decomposition. In Table 1, we display the results on some real and synthetic datasets used in Opolka et al. [5]. We observe that on all but one dataset (IMDB-BINARY), WT-GP-Hodge improves the classification accuracy. This may be surprising given that for vertex features, the Hodge decomposition only yields a co-exact and a harmonic part (since the exact part vanishes), and moreover, the harmonic part is just a constant value on the connected components of a graph. This implies that we can gain accuracy by splitting up vertex features into a constant bias (the harmonic part) and fluctuations around it (the co-exact part), and treating them separately with different kernels. However, when constant biases do not aid the classification, (e.g. when each class has vertex features of roughly equal magnitudes), then we do not expect WT-GP-Hodge to improve over WT-GP.
@@ 3.2 Vector field classification
In our second experiment, we consider the task of classifying noisy vector fields, i.e. predicting whether they are mostly divergence-free or curl-free. We proceed by generating 100 random vector fields, with half being mostly divergence-free (Figure 1(c)), and the other half mostly curl-free (Figure 1(d)). The generated vector fields are then projected onto the edges of a randomly generated triangular mesh on a square domain of $N$ vertices (Figure D.2). Finally, we corrupt the edge features with i.i.d. Gaussian noise. The resulting dataset consists of 100 oriented graphs, each containing scalar edge features corresponding to the net flow of the vector field along the edges. Again, we compare our method against the vanilla WT-GP classification method. However, since WT-GP does not take edge features, when using this, we first convert the graph to a line-graph before applying it. We thus refer to it as WT-GP-LG. In Table 2, we display the results of WT-GP-LG and WT-GP-Hodge for various choices of $N$. We see that WT-GP-Hodge is consistently better, with large improvements as mesh resolution is increased. On the other hand, WT-GP-LG cannot distinguish accurately between divergence-free and curl-free fields, even as the mesh resolution becomes higher. Likely reasons for this are: (1) the Hodge decomposition in WT-GP-Hodge can help to distinguish more clearly between divergence-free and curl-free features, and (2) there are properties that are canonical to edges, such as orientation, which WT-GP-Hodge can handle naturally, whereas WT-GP-LG cannot. In Figure 2, we also plot the classification accuracy with varying noise level, which further shows robustness of WT-GP-Hodge to noise compared to WT-GP-LG.
@@
Table 1: Comparison of classification accuracy on several graph classification benchmark datasets. Results display means and standard deviations across 10 seeds.
@@
Figure 1: Accuracy vs. noise level in the vector field classification on 10 seeds.
@@
Figure 2: Accuracy of vector field classification. Results are across 10 seeds.

### v2

@@ Abstract
The problem of classifying graphs is ubiquitous in machine learning. While it is standard to apply graph neural networks or graph kernel methods, Gaussian processes can be employed by transforming spatial features from the graph domain into spectral features in the Euclidean domain, and using them as input points. However, this approach only takes into account features on vertices, whereas some graph datasets also support features on edges. In this work, we present a Gaussian process-based classification algorithm that can leverage one or both vertex and edges features. Furthermore, we take advantage of the Hodge decomposition to better capture the intricate richness of vertex and edge features, which can be beneficial on diverse tasks.
@@ 1 Introduction
Classification is omnipresent in machine learning, yet typically assumes data to be Euclidean. Extending this task to non-Euclidean domains, such as graphs, presents challenges due to their irregularity, varying sizes, and multi-site information (e.g. vertices and edges). However, classifying graphs is of critical importance in scientific and industrial applications, being used for instance, to predict properties of molecules, or discovering new drugs [1, 2]. Although graph neural networks [3] are usually the model of choice for such applications, a downside is that they may require large datasets for effective training. Gaussian processes (GP) [4], on the other hand, prove to be a data-efficient and interpretable modelling choice. They do not need separate validation datasets to tune hyperparameters and provide robust uncertainty estimates for predictions. This makes them ideal for small-data scenarios and high-risk decision-making tasks that require reliable uncertainty estimates.
@@
In a recent work, Opolka et al. [5] introduced a GP-based algorithm capable of classifying graphs. Their method relies on tools developed from the graph signal processing literature [6], including the spectral graph Fourier transform [7] and the spectral graph wavelet transform [8]. Specifically, spectral graph methods use spatial graph features to compute graph spectral coefficients, which may be leveraged to generate spectral features in the Euclidean domain. Such spectral representations can be passed as input points to a standard GP [5], subsequently employed for classification via approximate inference [9]. Closely related are techniques developed in the early graph neural network literature, for instance, Spectral Networks [10] and ChebNet [11], which lean on the graph Fourier transform and graph wavelet transform to establish a notion of convolution on graphs. While the approach proposed in Opolka et al. [5] accommodates features living on vertices, it cannot easily take into account features on edges. However, edge information can often be as valuable as vertex information, representing crucial quantities such as flows [12, 13] and bond energies [14].
@@
Our paper aims to fill this gap by proposing a novel GP-based classification algorithm that naturally incorporate features on vertices, edges, and more generally, simplices, building on recent work defining GPs on simplicial complexes [15] and cellular complexes [16]. Moreover, we utilise the celebrated Hodge decomposition on spatial graph features, separating and processing them into three canonical components, each exhibiting distinct properties. This enables the modelling of these components separately, using different kernels. This idea have been used successfully in past works [17, 18, 19, 20, 15], providing greater flexibility. We are not aware that these recent techniques have been applied in the context of graph classification. We demonstrate empirically that these extensions achieve similar or better performance than the method introduced by Opolka et al. [5]. Although we focus on graphs, our approach can be extended easily to higher-order networks, such as simplicial complexes, cellular complexes, and hypergraphs, which can describe polyadic interactions [21, 22, 23, 24], thereby generalising the dyadic interactions found in graphs (see Appendix C). Finally, we highlight that our method can be readily adapted for regression by selecting an appropriate likelihood.
@@ 2 Gaussian Processes for Graph Classification
We assume a dataset containing $M$ undirected graphs $\mathcal{G}^{(1)},\dots,G^{(M)}$ and labels $y^{(i)}\in\mathbb{Z}$, for $1\leq i\leq M$. We assign an orientation to each graph but emphasise that this choice is arbitrary. Let $\mathcal{V}^{(i)}$ be the vertex set and $\mathcal{E}^{(i)}$ be the edge set, both finite. For each graph $\mathcal{G}^{(i)}=(\mathcal{V}^{(i)},\mathcal{E}^{(i)})$, there are $N_{v}^{(i)}$ vertices, $N_{e}^{(i)}$ edges, and an incidence matrix $\boldsymbol{B}_{ve}^{(i)}\in\mathbb{Z}^{N_{v}^{(i)}\times N_{e}^{(i)}}$. The latter encodes the incidence between each vertex and edge, and furthermore defines the graph Laplacian $\boldsymbol{L}_{v}^{(i)}\coloneq\boldsymbol{B}_{ve}^{(i)}\boldsymbol{B}_{ve}^{(i)\top}\in\mathbb{Z}^{N_{v}^{(i)}\times N_{v}^{(i)}}$. By considering $3$-cliques, we obtain the edge-to-triangle incidence matrix $\boldsymbol{B}_{et}^{(i)}\in\mathbb{Z}^{N_{e}^{(i)}\times N_{t}^{(i)}}$, where $N_{t}^{(i)}$ is the number of triangles in $\mathcal{G}^{(i)}$. Likewise, we define the graph Helmholtzian $\boldsymbol{L}_{e}^{(i)}\coloneq\boldsymbol{B}_{ve}^{(i)\top}\boldsymbol{B}_{ve}^{(i)}+\boldsymbol{B}_{et}^{(i)}\boldsymbol{B}_{et}^{(i)\top}\in\mathbb{Z}^{N_{e}^{(i)}\times N_{e}^{(i)}}$, which applies to edges rather than vertices. We underline that $\boldsymbol{L}_{v}^{(i)}$ and $\boldsymbol{L}_{e}^{(i)}$ are instances of the discrete Hodge Laplacian (see Appendix C.2). Let $\mathcal{V}^{(i)}\rightarrow\mathbb{R}^{D_{v}}$ and $\mathcal{E}^{(i)}\rightarrow\mathbb{R}^{D_{e}}$ be two functions, for $D_{v},D_{e}\in\mathbb{N}$. By introducing an ordering on vertices and edges, we represent the preceding functions as matrices $\mathbb{R}^{D_{v}\times N_{v}^{(i)}}$ and $\mathbb{R}^{D_{e}\times N_{e}^{(i)}}$, respectively, which are understood as vertex and edge feature matrices, containing $D_{v}$ and $D_{e}$ channels, respectively. We denote vertex features for channel $1\leq d\leq D_{v}$ by the column vector $\boldsymbol{x}^{(i)}_{vd}\in\mathbb{R}^{N_{v}^{(i)}}$ and edge features for channel $1\leq d\leq D_{e}$ by the column vector $\boldsymbol{x}^{(i)}_{ed}\in\mathbb{R}^{N_{e}^{(i)}}$. We stress that our approach can adapt to graphs that may have vertex features, edge features, or both.
@@ 2.1 Wavelet transforms on graphs
The key idea behind our graph classification algorithm is to convert vertex and edge features from the graph domain into spectral features in the Euclidean domain, enabling standard GP classification. We first consider the eigendecomposition of the graph Laplacian and graph Helmholtzian
@@
and define graph Fourier coefficients as projections of spatial graph features onto the eigenbases,
@@
We observe that $\hat{\boldsymbol{x}}^{(i)}_{\bullet d}$ reside in the eigenspace of $\boldsymbol{L}_{\bullet}^{(i)}$. Furthermore, they are invariant to vertex and edge ordering, making them sound choices for constructing spectral features, and they are perfectly localised in frequency, capturing global properties of the original features. However, it is often beneficial to also possess spatially localised information, focusing on local properties. A solution is to compute the more flexible graph wavelet coefficients [8] (see Appendix B) by modulating Fourier coefficients using a wavelet filter on the eigenvalues $\boldsymbol{\Lambda}_{v}^{(i)}$ and $\boldsymbol{\Lambda}_{e}^{(i)}$, and then perform the inverse Fourier transform. A wavelet filter is a combination of a scaling function at a single scale and a wavelet function at multiple scales, offering multi-scale resolution. Wavelet functions, $b:\mathbb{R}\rightarrow\mathbb{R}$ and $d:\mathbb{R}\rightarrow\mathbb{R}$, operate as band-pass filters, and scaling functions, $a:\mathbb{R}\rightarrow\mathbb{R}$ and $c:\mathbb{R}\rightarrow\mathbb{R}$, are low-pass filters. A wavelet filter captures one perspective of a graph, but to obtain a comprehensive picture, it is essential to employ multiple wavelet filters. Wavelet filters $j$ are defined by
@@
where $L_{\bullet}$ is the number of scales, $W_{\bullet}$ is the number of wavelet filters, and $\Theta_{\bullet j}$ is a collection of trainable parameters controlling the scaling. Wavelet coefficients are given by
@@ 2.2 Hodge decomposition
The discrete Hodge decomposition [25] (see Appendix C.3) states that spatial graph feature spaces, i.e. the spaces inhabited by $\boldsymbol{x}_{vd}^{(i)}$ and $\boldsymbol{x}_{ed}^{(i)}$, can each be separated into an orthogonal sum of three subspaces, exact, co-exact, and harmonic, collectively referred to as the Hodge subspaces. From this, the eigenbases in (1) can be divided into sub-eigenbases, each spanning a different Hodge subspace,
@@
We observe that $\boldsymbol{U}^{(i)}_{v}$ has only two components. The co-exact sub-eigenbasis $\boldsymbol{U}^{(i)}_{vc}$ amounts to the non-zero eigenvectors of $\boldsymbol{L}_{v}^{(i)}$, and the harmonic sub-eigenbasis $\boldsymbol{U}^{(i)}_{vh}$ to the zero ones. The exact and co-exact sub-eigenbases $\boldsymbol{U}^{(i)}_{ee}$ and $\boldsymbol{U}^{(i)}_{ec}$ are the non-zero eigenvectors of $\boldsymbol{B}_{ve}^{(i)\top}\boldsymbol{B}_{ve}^{(i)}$ and $\boldsymbol{B}_{et}^{(i)}\boldsymbol{B}_{et}^{(i)\top}$, respectively. Finally, the harmonic sub-eigenbasis $\boldsymbol{U}^{(i)}_{eh}$ comprises the zero eigenvectors of $\boldsymbol{L}_{e}^{(i)}$. For edges, the exact and co-exact components are sometimes termed gradient and curl components, respectively, reminiscent of vector fields. A gradient part is curl-free, indicating no vortices. A curl part is divergence-free, meaning no sources or sinks. A harmonic part is curl-free and divergence-free.
@@ 2.3 Hodgelet spectral features
We generate graph spectral features from graph wavelet coefficients and then use them in downstream GP classification tasks (see Appendix A). By combining the wavelet transform (5) and the Hodge decomposition (6), we compute the wavelet coefficients $\hat{\boldsymbol{x}}^{(i)}_{vdjc},\hat{\boldsymbol{x}}^{(i)}_{vdjh}$ and $\hat{\boldsymbol{x}}^{(i)}_{edje},\hat{\boldsymbol{x}}^{(i)}_{edjc},\hat{\boldsymbol{x}}^{(i)}_{edjh}$. We derive our Hodgelet spectral features by concatenating the 2-norm of the preceding wavelet coefficients across each wavelet filter and channel, resulting in the column vectors $\boldsymbol{v}^{(i)}_{c},\boldsymbol{v}^{(i)}_{h}\in\mathbb{R}^{W_{v}D_{v}}$ and $\boldsymbol{e}^{(i)}_{e},\boldsymbol{e}^{(i)}_{c},\boldsymbol{e}^{(i)}_{h}\in\mathbb{R}^{W_{e}D_{e}}$. The spectral representations are then fed to our additive Hodgelet kernel
@@
where $\kappa_{\bullet\sqbullet}$ is a standard kernel function, such as the squared exponential kernel function. We note that parameters $\Theta_{\bullet j}$ are optimised jointly with the kernel hyperparameters and that a separate kernel for each part of the Hodge decomposition offers greater flexibility. We highlight that our GP-based classification algorithm supports multi-dimensional spatial graph features and graphs of varying sizes, in contrast to typical graph kernel-based methods [26]. The GP component scales according to the number of graphs, while the eigendecompositions are a one-off cost that can be performed in advance.
@@ 3 Experiments
The aim of the experiments is two-fold: (1) we validate the added flexibility given by the Hodge decomposition, and (2) we demonstrate that when edge features are present, it is better to work with edges directly rather than converting graphs to line-graphs. The latter point has been observed in previous works [22, 15] and our experiments, across 10 seeds, further validate their conclusions.
@@ 3.1 Graph classification benchmarks
Our first experiment compares the method in Opolka et al. [5], which we refer to the wavelet-transform GP (WT-GP), which does not use the Hodge decomposition, to our method, WT-GP-Hodge, which employs the decomposition. In Table 1, we display the results on some standard graph classification benchmark datasets used in Opolka et al. [5]. We observe that on all but one dataset, WT-GP-Hodge improves the classification accuracy. This may be surprising as the Hodge decomposition for vertex features yields only co-exact and harmonic parts, where the harmonic part is constant across the connected components of the graph. This suggests that we can gain accuracy by separating vertex features into a constant bias (harmonic part) and fluctuations around it (co-exact part), using different kernels for each. However, if constant biases do not aid classification (e.g. when classes have vertex feature of similar magnitudes), then we do not expect WT-GP-Hodge to improve over WT-GP.
@@ 3.2 Vector field classification
Our second experiment consider the task of classifying noisy vector fields, i.e. predicting whether they are predominantly divergence-free or curl-free. We proceed by generating 100 random vector fields, with half mostly divergence-free (Figure 1(c)) and the other half mostly curl-free (Figure 1(d)). The generated vector fields are then projected onto the edges of a randomly generated triangular mesh on a square domain with $N$ vertices (Figure D.2). Finally, we corrupt the edge features with i.i.d. Gaussian noise, resulting in a dataset composed of $100$ oriented graphs, each containing scalar edge features corresponding to the net flow of the vector field along the edges. Again, we compare our method against the vanilla WT-GP classification method. However, since WT-GP does not take edge features, we first convert graphs to line-graphs before applying it. We refer to it as WT-GP-LG. In Table 2, we display the results of WT-GP-LG and WT-GP-Hodge for various choices of $N$. We see that WT-GP-Hodge is consistently better, with large improvements as mesh resolution is increased. On the other hand, WT-GP-LG cannot distinguish accurately between divergence-free and curl-free fields, even as the mesh resolution becomes higher. Likely reasons: (1) the Hodge decomposition in WT-GP-Hodge helps to discriminate more clearly between divergence-free and curl-free components, and (2) there are properties that are canonical to edges, such as orientation, which WT-GP-Hodge can handle naturally, whereas WT-GP-LG cannot. In Figure 2, we plot the classification accuracy with varying noise level, which shows robustness of WT-GP-Hodge to noise compared to WT-GP-LG.
@@
Table 1: Comparison of classification accuracy on several graph classification benchmark datasets.
@@
Figure 1: Accuracy vs. noise level in the vector field classification.
@@
Figure 2: Accuracy of vector field classification.